OpenStudy (anonymous):

Solve for x using properties of logarithms, 5^(x-1)=2^(2x+1)

OpenStudy (anonymous):

\[\log_{5/4} 10\]

OpenStudy (callisto):

\[5^{x-1}=2^{2x+1}\] Take log on both sides \[log5^{x-1}=log2^{2x+1}\]BY\[loga^{b}=bloga\] \[(x-1)log5=(2x+1)log2\] Can you solve it from here?

OpenStudy (anonymous):

I've gone that far, do I distribute?